What is the x and y intercepts of x-3y=9, y=7x+5, y=6x, and -4x+y=10
Answer by colliefan(242)

The x intercept is the point at which y=0 for an equation. The y intercept is the point at which x=0. Take any of these formulas and substitute 0 in for either the x or the y to determine the intercept.
For example, taking the first equation:
x-3y=9
If x=0 this equation becomes 0-3y=9. Solving for y gives y=-3. For this equation the y-intercept is -3.
For the same equation, if y=0 it becomes: x-3(0)=9, x=9. The x-intercept is 9.

The same thing can be done with the other equations.